OpenStudy (savannah_noelle):

How would I solve these equations using the substitution method? y=x-2 y=-x+8

OpenStudy (amistre64):

what does it mean to substitute?

OpenStudy (amistre64):

it means to replace with an equivalent form ..... if they insist that y=x-2. then every where you see a y, substitute it with the form: x-2, and solve for x

OpenStudy (savannah_noelle):

so the equation now becomes x-2=x-2?

OpenStudy (amistre64):

the top one does, which granted is of little use to us; but they give us that bottom one which helps out alot when we substitute :)

OpenStudy (amistre64):

y = -x+8 x - 2 = -x + 8 now that one we can get a good value for x from

OpenStudy (savannah_noelle):

How would I solve this?

OpenStudy (savannah_noelle):

would the orderedpair just be (-2,8)?

OpenStudy (anonymous):

Using substitution, we can substitute x-2 in place of y in the second equation. x - 2 = -x + 8. Add 2 to both sides. x = -x + 8 + 2. Add x to one side. x + x = 8 + 2. 2x = 10. Divide by 2. x = 5. So your first number in the ordered pair would be (5,y). Use the value found for x to find the value of y. y = x - 2. y = (5) - 2. Solve. :) This will give you the value for y in the ordered pair (x,y)

OpenStudy (savannah_noelle):

Thank you sooo much!

OpenStudy (amistre64):

since solving for 1 equation in 1 variable is covered extensively before solving 2 equations in 2 variables; i assumed you would have known how to find a solution to: x-2 = -x+8

OpenStudy (savannah_noelle):

Can you guys help me with one more? The question is x+2y=-3 x=-3y For x+2y=-3 I substituted in x and got -3y+2y=-3 From there I added a three to each side and got y+2=0 Then I subtracted 2 from each side and got y=-2 Then I substituted y into the equation x=-3y and I got 6 So in the end my ordered pair was (6,-2) The website is saying I got it wrong though, can you tell me what I did wrong?

OpenStudy (anonymous):

You mistake happened when you added three to each side. \[-3y + 2y \neq y\]

OpenStudy (anonymous):

Instead you should have gotten -y + 2 = 0.

OpenStudy (savannah_noelle):

I'm confused :/what do I do after I get -3y+2y=-3

OpenStudy (anonymous):

Oh, oops. -y + 2 = 0 is not right. You did good you just added the y's wrong :) And you want to make sure you keep the variable isolated, don't try to add numbers back onto the same side of the equation as the variable. -3y + 2y = -3. -1y = -3 y = 3.